OpenStudy (anonymous):

Write the equation y - 2 = 2(x-3) in standard form!

OpenStudy (anonymous):

First we can distribute on the right side: y-2=2x-6 Now we just need to do one more step. Which one?

OpenStudy (anonymous):

simplify it y-2=2x-6 since standard form is ax+by+c=0 so y+4=2x 2x-y-4=0

OpenStudy (anonymous):

@sami-21 We aren't supposed to bring answers, right?

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

@rafabc02 but we can post detail response. it is not just answer .it is with the steps .

OpenStudy (anonymous):

OK, great! Thanks!

OpenStudy (anonymous):

@Australian Do you have any doubt about the steps to get the answer? We are here to help you! ;)

OpenStudy (anonymous):

No, thank you so much for asking rafa.

OpenStudy (anonymous):

You are welcome! Don't forget to give the "best response" to sami-21 or to me. See you!